OpenStudy (anonymous):

Challenging Hyperbolic Proof:

OpenStudy (anonymous):

Prove that \[\frac{ 1+tanhx }{ 1-tanhx }=e ^{2x}\]

OpenStudy (anonymous):

Medal and fans for first. Already know it.

OpenStudy (anonymous):

I think it's just tedious algebra. Using the definition of tanh(x) = {e^x - e^(-x)] / (e^x + e^-x), substitute back into the equation and simplify

OpenStudy (anonymous):

You do not need (e^x-e^-x)/(e^x+e^-x) to solve at all. To prove it, it must be gone.

OpenStudy (anonymous):

I mean to prove it faster without tedious algebra, it should be gone.

OpenStudy (empty):

\[\frac{\cosh x }{\cosh x}\frac{ 1+\tanh x }{ 1-\tanh x }=\frac{\cosh x + \sinh x}{\cosh x - \sinh x} = \frac{e^x}{e^{-x}}\]

OpenStudy (anonymous):

You missed a between the first and second, but that seems reasonable.

OpenStudy (anonymous):

a step*

OpenStudy (empty):

What step? I skipped several steps but they were simple so no point in wasting time typing :P

OpenStudy (empty):

If you want me to elaborate on my reasoning anywhere I can explain myself. :D

OpenStudy (anonymous):



OpenStudy (anonymous):

Just to clarify. Its all good.

OpenStudy (empty):

I think the bigger jump is this one: \[\cosh x - \sinh x = e^{-x}\] since it's kind of subtle to see how this is true. \[\cosh x = \cosh -x\] and \[-\sinh x = \sinh -x\] so we can plug these in to get it.

OpenStudy (anonymous):

yea and that fact that sinhx+coshx=e^x while coshx-sinhx=e^-x